# Partial Search in a Frozen Network is Enough to Find a Strong Lottery Ticket

### v1

@@ Abstract
Randomly initialized dense networks contain subnetworks that achieve high accuracy without weight learning—strong lottery tickets (SLTs). Recently, Gadhikar et al. (2023) demonstrated theoretically and experimentally that SLTs can also be found within a randomly pruned source network, thus reducing the SLT search space. However, this limits the search to SLTs that are even sparser than the source, leading to worse accuracy due to unintentionally high sparsity. This paper proposes a method that reduces the SLT search space by an arbitrary ratio that is independent of the desired SLT sparsity. A random subset of the initial weights is excluded from the search space by freezing it—i.e., by either permanently pruning them or locking them as a fixed part of the SLT. Indeed, the SLT existence in such a reduced search space is theoretically guaranteed by our subset-sum approximation with randomly frozen variables. In addition to reducing search space, the random freezing pattern can also be exploited to reduce model size in inference. Furthermore, experimental results show that the proposed method finds SLTs with better accuracy and model size trade-off than the SLTs obtained from dense or randomly pruned source networks. In particular, the SLT found in a frozen graph neural network achieves higher accuracy than its weight trained counterpart while reducing model size by 40.3×40.3$\times$40.3 ×.
Deep Learning, Strong Lottery Ticket, Model Compression
@@
1 Introduction
The Strong Lottery Ticket Hypothesis (SLTH) conjectured the existence of subnetworks within a randomly weighted network—strong lottery tickets (SLTs)—that achieve comparable accuracy to trained dense networks (Zhou et al., 2019; Ramanujan et al., 2020; Malach et al., 2020). The existence of such subnetworks that do not require weight learning, illustrated in Figure 1(left), has been demonstrated experimentally (Zhou et al., 2019; Ramanujan et al., 2020; Huang et al., 2022; López García-Arias et al., 2023; Yeo et al., 2023) and proven theoretically for dense source networks (Malach et al., 2020; Orseau et al., 2020; Pensia et al., 2020; Diffenderfer & Kailkhura, 2021; da Cunha et al., 2021; Burkholz, 2022a; Burkholz et al., 2022; Ferbach et al., 2023; Gadhikar et al., 2023).
Since SLTs can be reconstructed from the random seed of the weights and a supermask (a simple binary pruning mask), potentially eliminating the need for storing and loading weights from memory, they offer a particularly advantageous opportunity for specialized inference hardware (Hirose et al., 2022). However, the existing algorithms for finding SLTs suffer from the same high computational cost as weight optimization, since they also employ backpropagation to update one parameter per neural connection (Ramanujan et al., 2020; Zhou et al., 2021; Diffenderfer & Kailkhura, 2021; Sreenivasan et al., 2022).
@@ 1 Introduction
Recently, Gadhikar et al. (2023) demonstrated that this steep training cost can be alleviated by randomly pruning the source network at initialization, reducing the number of edges to be optimized (see Figure 1, center). However, this limits the SLT search space to the high sparsity region, which may not offer the optimal accuracy.
@@
This paper introduces a novel method that reduces the search space by an arbitrary ratio independent of the desired SLT sparsity: in addition to random pruning at initialization, it also locks randomly chosen parameters at initialization to be a permanent part of the SLT (i.e., never pruned), as exemplified in Figure 1(right). Both the pruned and the locked parameters—the frozen parameters—are excluded from the search space, therefore reducing the training cost. Excluding them from the supermask also reduces the model size in inference, as the random freezing pattern can also be regenerated from its seed (Wimmer et al., 2020).
Moreover, our experiments show that SLTs found within frozen networks can achieve a better accuracy to model size trade-off than SLTs obtained from dense (Ramanujan et al., 2020; Sreenivasan et al., 2022) or sparse (Gadhikar et al., 2023) random networks. In addition to demonstrating this method experimentally, this paper also provides a mathematical analysis of the existence of SLTs within such partially frozen source networks.
In summary, this paper proposes a novel method that features the following contributions:
  • It reduces the SLT search space regardless of the desired SLT sparsity by random partial freezing (pruning and locking) of the source random network.
  • The proposed method is theoretically supported by our SLT existence result based on subset-sum approximation with randomly frozen variables.
@@
  • Experimental results show that SLTs found in the frozen network achieve better accuracy-to-search space and accuracy-to-model size trade-offs than SLTs found within dense or sparse random networks.
Figure 2: Burkholz (2022a) proved that target network of depth $L$ can be stochastically approximated by an SLT of depth $L\!+\!1$ by approximating each target weights ($w_{T}$) by a subset-sum of the source weights. We extend this proof to frozen source networks by solving the subset-sum approximation with some source parameters that are either unavailable ($w^{\prime}_{2}$) or must be used for the subset-sum ($w^{\prime}_{0}$).
@@ 2 Preliminaries
After outlining the background on strong lottery tickets (SLTs) and their search algorithms, this section recapitulates the subset-sum approximation technique for the existence proof of SLTs (Pensia et al., 2020; da Cunha et al., 2021; Burkholz, 2022b, a) and its extension to sparse random networks (Gadhikar et al., 2023), which Section 3 expands to the existence of SLTs within frozen networks.
@@ 2.1 Strong Lottery Tickets in Dense Networks
SLTs (Zhou et al., 2019; Ramanujan et al., 2020; Malach et al., 2020) are subnetworks within a randomly weighted neural network that achieve high accuracy without any weight training. Compared with learned dense weight models, SLTs can be reconstructed from a small amount of information: since the random weights can be regenerated from their seed, it is only necessary to store the binary supermask and the seed (Hirose et al., 2022). Instead of updating weights, SLT search algorithms for deep neural networks (Ramanujan et al., 2020; Zhou et al., 2021; Sreenivasan et al., 2022) use backpropagation to update weight scores, which are then used to generate the supermask. For example, the Edge-Popup algorithm (Ramanujan et al., 2020; Sreenivasan et al., 2022), used in this paper, finds SLTs by applying a supermask generated by from the connections with the top-$k\%$ scores. The need to keep scores in addition to the weights, and the sorting of scores necessary for generating the supermask make the computational cost of Edge-Popup slightly higher than conventional weight training.
@@ 2.2 SLT Existence via Subset-Sum Approximation
Given a set of random variables and a target value, the subset-sum problem consists on finding a subset whose total value approximates the target. Lueker (1998) showed that such subset exists with high probability if the number of random variables is sufficiently large:
Lemma 2.1 (Subset-Sum Approximation (Lueker, 1998)).
@@
Based on Lemma 2.1, previous works (Pensia et al., 2020; da Cunha et al., 2021; Burkholz, 2022b, a) showed that an SLT that approximates an arbitrary target network exists in a dense source network if it is wider and deeper than the target network. For example, Burkholz (2022a) proved that a source network with depth $L\!+\!1$ and larger width than the target network contains an SLT that can approximate the target network with depth $L$. As exemplified in Figure 2 (left), to approximate a weight $w_{T}$ of the target network, the SLT uses weights $w_{0},w_{1}$, and $w_{2}$ from the source network to construct the subset-sum. Therefore, the source network is required to be sufficiently wider than the target network, as stated in Lemma 2.1.
@@ 2.3 Strong Lottery Tickets in Sparse Networks
Recently, Gadhikar et al. (2023) revealed that SLTs also exist within sparse source networks, i.e., random networks that have been randomly pruned at initialization (see Figure 1, center). In addition to their theoretical contributions, their approach offered a practical advantage: since there are less parameters available from the beginning, the search space is reduced in proportion to the initial pruning ratio. This search space reduction can be exploited for reducing the computational cost of training.
@@ 2.4 SLT Existence in Sparse Networks
To prove the existence of SLTs within sparse networks, Gadhikar et al. (2023) extended the subset-sum approximation (Lemma 2.1) to the situation where randomly chosen variables are permanently pruned at initialization:
Lemma 2.2 (Subset-Sum Approximation in Sparse Networks (Gadhikar et al., 2023)).
@@
By applying this extended lemma instead of Lemma 2.1 to the SLT existence theorem presented by Burkholz (2022a), Gadhikar et al. (2023) proved the SLT existence in sparse networks as the following theorem:
Theorem 2.3 (SLT Existence in Sparse Networks (Gadhikar et al., 2023)).
Let a target network $f_{T}$ with depth $L$ and a sparse source network $f_{S}$ with depth $L\!+\!1$. Assume that the source network is randomly pruned with pruning ratio $p_{l}$ for each $l$-th layer. Also assume that these networks have ReLU activation function and initialized by using a uniform distribution $U[-1,1]$. Then, except with exponentially small probability, a subnetwork $f_{SLT}$ exists in the sparse source network $f_{S}$ such that $f_{SLT}$ approximates the target network $f_{T}$ if the width of $f_{S}$ is sufficiently large.
@@ 3 Strong Lottery Tickets in Frozen Networks
Although pruning at initialization provides a search space reduction, it also imposes a limitation on the sparsity range, as it can only search for SLTs with a sparsity higher than the initial pruning ratio. For example, although an initial pruning ratio of $90\%$ vastly reduces the search space, it limits the search to only SLTs sparser than $90\%$. Therefore, this reduction in search space is incompatible with exploring the whole sparsity range for optimal accuracy.
This paper proposes a novel method that reduces the search space regardless of the desired SLT sparsity to capitalize on the training cost reduction without sacrificing accuracy. We explore the existence and performance of SLTs within a frozen source network, i.e., a random network that, in addition to randomly pruned parameters, has randomly chosen parameters forced to be permanent (i.e., never pruned) part of the SLT—locked parameters (see Figure 1, right). Since freezing is performed at initialization, before the SLT search, frozen parameters do not have a score and are not eligible for exploration, thus reducing the search space. Locking allows to further reduce the parameter search without limiting the initial pruning ratio, allowing to reduce the search space even in the case of dense source networks.
This section first provides a theoretical analysis on the existence of SLTs in such a frozen network by developing a new extension of the subset-sum approximation introduced in Section 2. Then, based on this analysis, it proposes partial SLT search, a novel method for finding such SLTs with a reduced search space. In preparation for the evaluation experiments in Section 4, we also perform a preliminary investigation of the optimal settings of the proposed method.
@@ 3.1 SLT Existence in Frozen Networks
Here we provide a theoretical result that indicates that an SLT capable of approximating a target network exist in a frozen network. This is proved by extending the subset-sum approximation lemma (Lemma 2.1) to the case where some parameters are locked (for detailed proof see Appendix A).
@@
Then, combining Lemma 3.1 with Lemma 2.2 further extends the subset-sum approximation to the situation where some random variables are frozen (i.e., pruned or locked).
@@
Finally, by applying Lemma 3.2 to the Theorem 2.5 in Gadhikar et al. (2023) instead of Lemma 2.2, it follows that an SLT, which approximates a target network, exists in frozen networks.
@@
Figure 4: Comparison of different prune:lock proportions of a $80\%$ freezing ratio using a Conv6 on CIFAR-1010{10}10.
@@ 3.2 Partial SLT Search in Frozen Networks
Figure 4 illustrates how network freezing reduces the search space. The total amount of frozen parameters is determined by a global freezing ratio $F_{r}$, which is the sum of the respective pruning ratio $P_{r}$ and locking ratio $L_{r}$. Therefore, the SLT search space $S$ is formed by the $1\!-\!F_{r}$ non-frozen parameters of the original dense network. As explained previously, it is not possible to search for an SLT with a lower sparsity (denser) than the source network. Consequently, the pruning ratio sets the lower bound of the search space. On the other hand, it is not possible to search for an SLT that prunes more parameters than available, so the locking ratio sets the upper bound of the search space. (Specifically, the upper bound is given by $1\!-\!F_{r}$) Thus, these ratios allow to freely control the size and position of the search space over the SLT sparsity spectrum.
@@
  • Edge Per Layer (EPL): Each layer’s pruning ratio is set so that they all have the same number of remaining weights (Price & Tanner, 2021; Gadhikar et al., 2023).
  • Erdős-Rényi-Kernel (ERK): The pruning ratio of layer $l$ is proportional to the scale $(C_{in}^{(l)}+C_{out}^{(l)}+k_{h}^{(l)}+k_{w}^{(l)})/(C_{in}^{(l)}\times C_{out}^{(l)}\times k_{h}^{(l)}\times k_{w}^{(l)})$, where $C_{in}^{(l)}$, $C_{out}^{(l)}$, $k_{h}^{(l)}$, and $k_{w}^{(l)}$ denote input channels, output channels, kernel height, and kernel width of the layer $l$, respectively (Evci et al., 2020).
The same strategy is employed to determine the layer-wise pruning and freezing ratios from their respective global ratios, and then locking ratios are calculated as the difference between the corresponding freezing and pruning ratios.
Freezing mask compression:
The freezing pattern can be encoded during inference as a ternary mask—a freezing mask—that indicates whether a parameter is pruned, locked, or part of the supermask. For example, encoding pruning as $-1$, locking as $+1$, and supermask inclusion as 00, the layer-wise freezing mask can be encoded as $\bm{m}^{(l)}_{l}+(\bm{m}^{(l)}_{p}-1)$. Since this freezing mask is also random and fixed, it can be regenerated from its seed and ratios, similarly to the random weights. Furthermore, the supermask size is reduced by excluding from it the frozen parameters, so the SLTs found by this method can be compressed in inference to an even smaller memory size than those produced by the existing SLT literature (Hirose et al., 2022; Okoshi et al., 2022; López García-Arias et al., 2023).

### v2

@@ Abstract
Randomly initialized dense networks contain subnetworks that achieve high accuracy without weight learning—strong lottery tickets (SLTs). Recently, Gadhikar et al. [10] demonstrated that SLTs can also be found within a randomly pruned source network, thus reducing the SLT search space. However, this limits the search to SLTs that are even sparser than the source, leading to worse accuracy due to unintentionally high sparsity. This paper proposes a method that reduces the SLT search space by an arbitrary ratio independent of the desired SLT sparsity. A random subset of the initial weights is excluded from the search space by freezing it—i.e., by either permanently pruning them or locking them as a fixed part of the SLT. In addition to reducing search space, the proposed random freezing can also provide the benefit of reducing the model size for inference. Furthermore, experimental results show that the proposed method finds SLTs with better accuracy-to-model size trade-off than the SLTs obtained from dense or randomly pruned source networks. In particular, the SLTs found in Frozen ResNets on image classification using ImageNet significantly improve the accuracy-to-search space and accuracy-to-model size trade-offs over SLTs within dense (non-freezing) or sparse (non-locking) random networks.
1 Introduction
The Strong Lottery Ticket Hypothesis (SLTH) conjectured the existence of subnetworks within a randomly weighted network—strong lottery tickets (SLTs)—that achieve comparable accuracy to trained dense networks [41, 30, 25]. The existence of such subnetworks that do not require weight training, illustrated in Figure 2(left), has been demonstrated experimentally [41, 30, 17, 21, 39] and proven theoretically for dense source networks [25, 27, 28, 6, 5, 1, 3, 8, 10].
SLTs offer a particularly advantageous opportunity for specialized inference hardware [14] since the random weights can be reconstructed from the seed (i.e., they do not need to be memorized), and the sparse binary pruning masks—supermasks—can be greatly compressed with entropy coding, vastly reducing off-chip memory access and its associated overheads. Furthermore, the binary nature of both the supermasks and the random weights can be exploited for multiplier-less execution, as demonstrated practically by Hirose et al. [14]. On the other hand, the existing training algorithms for finding SLTs suffer from the same high computational cost as weight optimization, since they also employ backpropagation to update one parameter per neural connection [30, 42, 6, 33].
@@ 1 Introduction
Figure 2: Freezing ($\bullet$) improves the accuracy-to-search space trade-off over pruning only (non-locking) ($\bullet$) or non-freezing ($\bullet$).
@@
Recently, Gadhikar et al. [10] demonstrated that this steep training cost can be alleviated by randomly pruning the source network at initialization, reducing the number of edges to be optimized (see Figure 2, center). However, by its nature, this method can only search SLTs in relatively high sparsity region, which also includes highly sparse networks with terrible accuracy such as layer collapse [11, 34]. Therefore, it would be desirable if we could reduce the search space while avoiding too sparse networks.
@@
This paper introduces a novel method that reduces the search space by an arbitrary ratio independent of the desired SLT sparsity: in addition to random pruning at initialization, it also locks randomly chosen parameters at initialization to be a permanent part of the SLT (i.e., never pruned), as exemplified in  Figure 2(right). Both the pruned and the locked parameters—the frozen parameters—are excluded from the search space, therefore reducing the training cost. Far from negatively impacting performance, this cost reduction proves efficient: the SLTs found within this partial search space are more accurate than those resulting from conventional methods, as shown in Figure 2. Moreover, resulting from the fact that this method can also be exploited to reduce model size in inference, our experiments show that SLTs found within frozen networks can achieve a better accuracy-to-model size trade-off than SLTs obtained from dense [30, 33] or sparse [10] random networks.
The contributions of this paper can be summarized as follows:
  • We propose a novel method that reduces the SLT search space with any desired sparsity by random partial freezing (pruning and locking) of the source random network. The soundness of our method is theoretically supported by our existence theorem of SLTs in such a randomly reduced search space, based on subset-sum approximation with randomly frozen variables.
  • We experimentally validated our method in three scenarios with highly accurate SLTs in low, medium, and high SLT sparsity regions, which reveal that parameter freezing consistently reduces the size of search spaces. Even with randomly frozen parameters, our method finds highly accurate SLTs that cannot be found within dense networks in some desired sparsities.
@@
  • Furthermore, the experimental results show that SLTs found in the frozen network achieve better accuracy-to-search space and accuracy-to-model size trade-offs than SLTs found within dense (non-freezing) or sparse (non-locking) random networks.
As mentioned above, SLTs are quite attractive for neural engine design, as they can vastly reduce the memory size for model storage, meaning that off-chip memory access—by far the major source of energy and time consumption [15]—can be drastically reduced for energy-efficient inference acceleration [14]. These contributions could potentially lead to attracting attention to SLTs also as an approach for efficient training, acquiring high-quality subnetworks with reduced training costs.
@@ 2 Preliminaries
After outlining the background on strong lottery tickets (SLTs) and their search algorithms, this section recapitulates the subset-sum approximation technique for the existence proof of SLTs [28, 5, 2, 1] and its extension to sparse random networks [10], which Section 3 expands to the existence of SLTs within frozen networks.
@@ 2.1 Strong Lottery Tickets in Dense Networks
SLTs [41, 30, 25] are subnetworks within a randomly weighted neural network that achieve high accuracy without any weight training. Compared with learned dense weight models, SLTs can be reconstructed from a small amount of information: since the random weights can be regenerated from their seed, it is only necessary to store the binary supermask and the seed [14]. Instead of updating weights, SLT search algorithms for deep neural networks [30, 42, 33] use backpropagation to update weight scores, which are then used to generate the supermask. For example, the Edge-Popup algorithm [30, 33] used in this paper finds SLTs by applying a supermask generated from the connections with the top-$k\%$ scores. The need to keep scores in addition to the weights and the score sorting operation necessary for generating the supermask make the computational cost of Edge-Popup slightly higher than conventional weight training.
Figure 3: We extend the SLT existence proof by Burkholz [1] (left) to frozen source networks (right) by solving the subset-sum approximation with some source parameters that are either unavailable ($w^{\prime}_{2}$) or must be used for the subset-sum ($w^{\prime}_{0}$).
@@ 2.2 SLT Existence via Subset-Sum Approximation
Given a set of random variables and a target value, the subset-sum problem consists on finding a subset whose total value approximates the target. Lueker [24] showed that such subset exists with high probability if the number of random variables is sufficiently large:
Lemma 2.1 (Subset-Sum Approximation [24]).
@@
Based on Lemma 2.1, which was first introduced into the SLT context by Pensia et al. [28], previous works [1, 2, 5, 28] showed that an SLT that approximates an arbitrary target network exists in a dense source network if it is logarithmically wider and constantly deeper than the target network. In particular, Burkholz [1] proved that a source network with depth $L\!+\!1$ and larger width than the target network contains an SLT that can approximate the target network with depth $L$. As exemplified in Figure 3 (left), to approximate a weight $w_{T}$ of the target network, the SLT uses weights $w_{0},w_{1}$, and $w_{2}$ from the source network to construct the subset-sum. Therefore, the source network is required to be sufficiently wider than the target network, as stated in Lemma 2.1.
@@ 2.3 Strong Lottery Tickets in Sparse Networks
Recently, Gadhikar et al. [10] revealed that SLTs also exist within sparse source networks, i.e., random networks that have been randomly pruned at initialization (see Figure 2, center). In addition to their theoretical contributions, their approach offered a practical advantage: since there are less parameters available from the beginning, the search space is reduced in proportion to the initial pruning ratio. This search space reduction can be exploited to reduce the computational cost of training.
@@ 2.4 SLT Existence in Sparse Networks
To prove the existence of SLTs within sparse networks, Gadhikar et al. [10] extended the subset-sum approximation (Lemma 2.1) to the situation where randomly chosen variables are permanently pruned at initialization:
Lemma 2.2 (Subset-Sum Approximation in Sparse Networks [10]).
@@
By applying this extended lemma instead of Lemma 2.1 to the SLT existence theorem presented by Burkholz [1], Gadhikar et al. [10] proved the SLT existence in sparse networks as the following theorem:
Theorem 2.3 (SLT Existence in Sparse Networks [10]).
Let a target network $f_{T}$ with depth $L$ and a sparse source network $f_{S}$ with depth $L\!+\!1$ be given. Assume that the source network is randomly pruned with pruning ratio $p_{l}$ for each $l$-th layer. Also assume that these networks have ReLU activation function and initialized by using a uniform distribution $U[-1,1]$. Then, except with exponentially small probability, a subnetwork $f_{SLT}$ exists in the sparse source network $f_{S}$ such that $f_{SLT}$ approximates the target network $f_{T}$ if the width of $f_{S}$ is sufficiently large.
@@ 3 Strong Lottery Tickets in Frozen Networks
Although pruning at initialization provides a search space reduction, it also imposes a limitation on the sparsity range, as it can only search for SLTs with a sparsity higher than the initial pruning ratio. For example, an initial pruning ratio of $90\%$ vastly reduces the search space, but it limits the search to only SLTs sparser than $90\%$. Therefore, this reduction in search space is incompatible with exploring the whole sparsity range for optimal accuracy.
This paper proposes a novel method that reduces the search space regardless of the desired SLT sparsity to capitalize on the training cost reduction without sacrificing accuracy. We explore the existence and performance of SLTs within a frozen source network, i.e., a random network that (in addition to randomly pruned parameters) has randomly chosen parameters forced to be permanent part (i.e., never pruned) of the SLT—locked parameters (see Figure 2, right). Since freezing is performed at initialization before the SLT search, frozen parameters do not have a score and are not eligible for exploration, thus reducing the search space. Locking allows to further reduce the parameter search without limiting the initial pruning ratio, reducing the search space even in the case of dense source networks.
This section first provides a theoretical analysis of the SLT existence in such frozen networks by developing a new extension of the subset-sum approximation introduced in Section 2. Then, based on this analysis, it proposes partial SLT search, a novel method for finding such SLTs with a reduced search space. In preparation for the evaluation experiments in Section 4, we also perform a preliminary investigation of the optimal settings of the proposed method.
@@ 3.1 SLT Existence in Frozen Networks
Here we provide a theoretical result indicating that an SLT capable of approximating a target network exists in a frozen network. This result is proved by extending the subset-sum approximation lemma (Lemma 2.1) to the case where some parameters are locked (for detailed proof see Appendix A).
@@
Then, by combining Lemma 3.1 and Lemma 2.2, we extend the subset-sum approximation to the situation where some random variables are frozen (i.e., pruned or locked).
@@
Finally, by applying Lemma 3.2 to the Theorem 2.5 in Gadhikar et al. [10] instead of Lemma 2.2, it follows that an SLT approximating a target network exists within a frozen network.
@@ 3.2 Partial SLT Search in Frozen Networks
Figure 4 illustrates how network freezing reduces the search space. The total amount of frozen parameters is determined by a global freezing ratio $F_{r}$, which is the sum of the respective pruning ratio $P_{r}$ and locking ratio $L_{r}$. Therefore, the SLT search space $S$ is formed by the $1\!-\!F_{r}$ non-frozen parameters of the original dense network. As explained previously, it is not possible to search for an SLT with a lower sparsity (denser) than the source network. Consequently, the pruning ratio sets the lower bound of the search space. On the other hand, it is not possible to search for an SLT that prunes more parameters than those available, so the locking ratio sets the upper bound of the search space. (Specifically, the upper bound is given by $1\!-\!L_{r}$.) Thus, these ratios allow to freely control the size and position of the search space over the SLT sparsity spectrum.
@@
  • Edge Per Layer (EPL): Each layer’s pruning ratio is set so that they all have the same number of remaining weights [29, 10].
  • Erdős-Rényi-Kernel (ERK): The pruning ratio of layer $l$ is proportional to the scale $(C_{in}^{(l)}+C_{out}^{(l)}+k_{h}^{(l)}+k_{w}^{(l)})/(C_{in}^{(l)}\times C_{out}^{(l)}\times k_{h}^{(l)}\times k_{w}^{(l)})$, where $C_{in}^{(l)}$, $C_{out}^{(l)}$, $k_{h}^{(l)}$, and $k_{w}^{(l)}$ denote input channels, output channels, kernel height, and kernel width of the layer $l$, respectively [7].
The same strategy is employed to determine the layer-wise pruning and freezing ratios from their respective global ratios, and then the locking ratios are calculated as the difference between the corresponding freezing and pruning ratios.
Freezing pattern encoding for model compression:
The freezing pattern can be encoded during inference as a ternary mask—a freezing mask—that indicates whether a parameter is pruned, locked, or part of the supermask. For example, encoding pruning as $-1$, locking as $+1$, and supermask inclusion as 00, the layer-wise freezing mask can be encoded as $\bm{m}^{(l)}_{l}+(\bm{m}^{(l)}_{p}-1)\in\{-1,0,1\}^{n^{(l)}}$. Since this freezing mask is also random and fixed, it can be regenerated from its seed and ratios, similarly to the random weights. Furthermore, the supermask size is reduced by excluding from it the frozen parameters, so the SLTs found by this method can be compressed in inference to an even smaller memory size than those produced by the existing SLT literature [14, 26, 21]. Also, this memory size compression reduces costly off-chip memory access on specialized neural inference accelerators [14, 4], thus SLTs within frozen networks have the potential to perform even faster and more energy-efficient inference processing.